# FoAM: Foresight-Augmented Multi-Task Imitation Policy for Robotic Manipulation

### v1

@@ Abstract
Multi-task imitation learning (MTIL) has shown significant potential in robotic manipulation by enabling agents to perform various tasks using a unified policy. This simplifies the policy deployment and enhances the agent’s adaptability across different contexts. However, key challenges remain, such as maintaining action reliability (e.g., avoiding abnormal action sequences that deviate from nominal task trajectories), distinguishing between similar tasks, and generalizing to unseen scenarios. To address these challenges, we introduce the Foresight-Augmented Manipulation Policy (FoAM), an innovative MTIL framework. FoAM not only learns to mimic expert actions but also predicts the visual outcomes of those actions to enhance decision-making. Additionally, it integrates multi-modal goal inputs, such as visual and language prompts, overcoming the limitations of single-conditioned policies. We evaluated FoAM across over 100 tasks in both simulation and real-world settings, demonstrating that it significantly improves IL policy performance, outperforming current state-of-the-art IL baselines by up to 41% in success rate. Furthermore, we released a simulation benchmark for robotic manipulation, featuring 10 task suites and over 80 challenging tasks designed for multi-task policy training and evaluation. See project homepage https://projFoAM.github.io/ for project details.
@@ I Introduction
One of the main goals in robot learning is to develop a general-purpose agent capable of performing various tasks based on user commands. For manipulation tasks, multi-task imitation learning (MTIL) serves as a key approach that enables agents to learn different tasks from expert demonstrations, eliminating the need for complex, hard-coded solutions or reward functions and leading to efficient and generalizable policies. However, in practical settings, achieving both generalization and reliability in MTIL remains a key challenge [1]. The agent must develop task-agnostic skills to generalize to new tasks and environments, while also capturing task-specific details to ensure reliable execution for individual tasks [2, 3].
@@
Previous research has shown that task adaptation in MTIL can be achieved by incorporating goal conditions into multi-task policy training [4, 5, 6, 7, 8, 9, 10, 11]. However, the low reliability of multi-task policies remains a persistent challenge. Existing MTIL policies that align robotic actions with expert actions based on goal conditions often fail to reason about these ambiguities and variations in expert demonstration data, severely impacting the agents’ performance on individual tasks. Meanwhile, single-goal conditioned policies [10, 12, 13, 14, 15, 16, 17, 9] come with its own limitations. For instance, policies conditioned on language instructions struggle to generalize to unseen tasks without data augmentation. While policies conditioned on goal images offer fine-grained guidance, they frequently encounter ambiguities in task activation and necessitate human intervention to accurately acquire and interpret the goal images. For instance, when the robot was tasked with placing an object into a multi-compartment locker, the resulting goal image (an object that was in the initial image disappeared in the goal image) was ambiguous due to occlusion, making it impossible to determine which specific compartment the object was placed in (see Figure 4).
@@
In this paper, we introduce the Foresight-Augmented Manipulation Policy (FoAM), a novel multi-task imitation learning policy designed to generate fine-grained actions while considering their consequences. This approach is inspired by how humans perform tasks, refining actions by comparing expected outcomes with the intended goal until the desired result is achieved [18]. Similarly, FoAM processes observation inputs, task prompts, and goal images to generate robotic actions and predict embeddings that represent the outcomes of those actions. During training, we apply an action loss to refine the policy’s behavior and introduce a foresight loss to ensure the policy accounts for the consequences of its actions. This foresight allows the agent to reason about its action across diverse tasks, and handle the ambiguities and variations in expert demonstration data, leading to more forward-looking and precise action. Additionally, FoAM leverages a fine-tuned vision-language model (VLM) [19] to autonomously generate goal images, enhancing the agent’s autonomy and improving its ability to generalize to unseen tasks and scenarios. Our approach demonstrates significant effectiveness, with evaluations across more than 100 tasks in both simulation and real world. FoAM outperforms state-of-the-art baselines, achieving an increase in success rate by up to 41% in success rate. Our main contributions are summarized as follows:
  1. 1.
    We introduce the Foresight-Augmented Manipulation Policy (FoAM), a novel multi-task IL policy with the ability to reason about the consequences of its actions.
  2. 2.
    We integrate a Vision-Language Model (VLM) as the Goal Imagination Module for FoAM, fusing language prompts with goal images, enhancing its ability to generalize to unseen tasks and scenarios.
  3. 3.
    We opensource a simulation benchmark with 10 task suites and over 80 tasks, alongside a real-world dataset comprising 14 tasks. We demonstrate that FoAM achieves state-of-the-art performance on these tasks.
@@ II Related Work
Goal-Guided Learning for Robotic Manipulation. In recent years, significant progress has been made in single-task learning policies [20, 21, 22, 23, 24]. However, to enable a wider adoption, intelligent robots must be equipped with the ability to adapt to diverse tasks and complete them effectively. Among current MTIL approaches, language-guided manipulation policies utilize large-scale datasets to achieve task generalization, or apply data augmentation techniques, such as vision generation models, to modify backgrounds and manipulated objects, enabling generalization across more tasks and scenarios with limited data [6, 25, 7, 9, 10, 26, 27, 8, 28, 29, 30, 31]. Despite the promising initial success, we found that language-conditioned policies often struggle with unseen tasks without sufficient data or extensive additional data augmentation. In parallel, other approaches have introduced goal images as task conditions [12, 10, 13, 32, 33, 34]. Compared to language inputs, images encapsulate richer information, enabling stronger generalization capabilities, and even allowing agents to perform zero-shot tasks [13]. However, goal images are susceptible to task ambiguity, where visually identical outcomes can be produced by different tasks, causing incorrect task activation.
@@
Moreover, the need to collect goal images from humans reduces agent autonomy. To leverage the strengths of both images and language, recent work has explored multi-modal prompts [15, 14, 35, 36, 34]. These methods extract entities from both language and image prompts, then compose a multi-modal prompt embedding based on predefined templates. In contrast, our approach leverages vision-language models to directly generate a goal image with high-dimensional semantic information. This generated goal image, along with the task prompt, is processed to further infer actions and predict future states. By doing so, our method addresses the limitations of goal-conditioned policies that rely solely on language prompts, goal images, or predefined templates.
@@
Agents with Vision Language Models. In recent years, Vision-Language Models (VLMs) have been introduced to robotics [30, 37, 38, 39, 40, 41, 36, 42, 43], enabling more complex visual reasoning and multimodal tasks. Meanwhile, in the community of image editing, VLMs also demonstrated the ability to understand language prompts, edit real-world images and produce highly realistic visual effects[44, 33]. [45, 14, 33, 46, 47] further validates that edited images generated via VLMs can be interpreted by robotic agents, where the generated goal images are directly used as a task activation condition. In contrast, we integrate VLMs seamlessly into our framework not only during the inference stage, but also in model training. Specifically, the VLM-generated goal images serve as the “labels” to compute the reconstruction loss for the foresight augmentation module (Section III-C), coupling the action policy learning and action result prediction.
@@ III METHOD
We seek to develop an imitation learning policy with strong generalization and robustness, empowering the agent to effectively complete a variety of tasks. To achieve this, we introduce FoAM, a novel multi-modal goal-conditioned policy learning framework. In the following sections, we will provide an overview of the FoAM in III-A, detail the fine-tuning of a cutting-edge visual-language model to generate goal images for FoAM in III-B, propose the foresight-augmented module in III-C, and introduce the implementation details of FoAM in III-D.
Figure 1: Training and inference pipelines of FoAM. The inputs remain the same during both training and inference. During training, actions and their consequences are predicted and compared with ground truth to update FoAM’s parameters. During inference, the trained policy predicts the action $\hat{a}_{t}$.
@@ III-A Pipeline Overview
The pipeline of FoAM is illustrated in Figure 1. FoAM is a transformer-based [48] policy that inherits the architecture of prior works [20] and is trained as a conditional variational autoencoder (CVAE) [49, 50]. The process begins with the user providing a task prompt $g_{l}$ to the agent based on the current observation, which is then fed into FoAM as a language goal condition through a pre-trained text encoder [51]. Concurrently, the task prompt, such as Put eggplant into the green bowl, is input into the Goal Imagination Module along with the initial observation $o_{i}$. The generated imaginary goal image or human collected real goal image $g_{i}$ serves as an additional condition. By integrating the agent’s proprioception $j_{t}$, visual observations $o_{t}$, and the latent style variable $z$ encoded from the action and joint data, the multi-goal conditional policy $\pi_{\theta}(\hat{a}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$ is learned through gradient descent during training. During inference, action chunks are predicted and robust actions are produced through temporal aggregation [20].
@@ III-B Fine-tuned Goal Imagination Module
We chose InstructPix2Pix (Ip2p)[19] as our goal imagination module, utilizing approximately 20,000 pairs of training data. Of these, 16,000 pairs were derived from the cleaning robot expert demonstrations provided by RT-1[25, 6, 7]. In this process, the first and last frames of the demonstrations were used as the original and edited images, respectively, with the corresponding task name serving as the instruction. Given that many of the demonstration datasets contained perturbations in the final frames caused by robot arm movements, we undertook a detailed data cleaning procedure to remove noise and ensure the training data quality. Additionally, we incorporated over 4,000 data pairs sourced from our own simulation and real-world datasets (see Section IV-A for details). We fine-tuned the model for 500 epochs on a single NVIDIA H100 GPU, a process that required approximately 3 days. During the image generation stage, with the model weights pre-loaded, processing each initial observation of size $480\times 640\times 3$ took about 4 seconds. Figure 2 presents a selection of initial observations alongside their corresponding imagined goal images, captured during VLM-FoAM joint inference experiments (see Section IV-C for details) in real-world. These results highlight the model’s ability to generate realistic and contextually visual effects based on the given initial observations and task prompts. We anticipate that these imagined visual scenes will help address the limitations of single-goal conditional policies, and allow the robot autonomously determine goal conditions, improving the robot’s generalization performance.
@@
Figure 2: Inference demonstration of the fine-tuned goal imagination module. The leftmost image illustrates the initial observation, while the next four images represent the imagined goal images generated based on the initial observation and the task prompt provided at the top. Please visit the project homepage for more examples.
III-C Foresight-augmented Module
@@
Humans possess exceptional cognitive abilities when it comes to understanding and interacting with external objects and events. This cognitive prowess enables individuals to not only perceive the current environment but also to anticipate future outcomes based on their actions. When tasked with a specific objective, humans often mentally envision the goal image even before initiating the task. Inspired by this human capability, we developed a foresight-augmented (FA) module to endow agents with a similar foresight mechanism. The FA enables the agent to concurrently comprehend both the actions it takes and the subsequent outcomes these actions will produce. By integrating this foresight into the agent’s decision-making process, we significantly improve the agent’s overall manipulation task performance.
@@
We train FoAM as a CVAE, utilizing an encoder similar to those in [20, 9, 52, 22] to generate the latent variable $z\sim q_{\phi}(z|a_{t:t+k},{j}_{t})$. The decoder is defined as policy $\pi_{\theta}(\hat{a}_{t:t+k},\hat{f}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$, which predicts a $k\times n$-dimensional action chunk $\hat{a}_{t:t+k}$ and foresight sequence $\hat{f}_{t:t+k}$ based on current observation and conditions, where $k$ represents the hyperparameter chunk size and $n$ denotes the dimension of the agent’s controlled action space. To enhance the agent’s ability to interpret and react to extended temporal contexts, we strategically increase the value of $k$, thereby broadening the agent’s vision of foresight. This adjustment is particularly important in the context of the multi-task dataset, where the length of each expert demonstration varies. For shorter tasks, we extend the data by replicating the final frame up to the maximum time step $T$ of all demonstrations, setting $k$ to be close to $T$. As the action chunk is being predicted, the FA module envisions $k$ potential foresight scenes, and selects the frame $\hat{g}_{i}=\hat{f}_{t:t+k}[k-t]$ that is temporally aligned with the goal image. This selected frame is then compared against the goal image $g_{i}$ to compute the foresight loss $L_{\text{foresight}}$. This process successfully emulates the strong cognitive abilities humans exhibit when approaching tasks, and we demonstrated in experiments that the FA module significantly enhances the agent’s task performance.
III-D FoAM Policy Implementation
FoAM is designed as a transformer-based policy with sufficient capacity to predict specific sequences by effectively integrating sequence information from the input. FoAM is implemented using an ACT-like architecture [20] with the CVAE framework. The language conditional embedding is obtained by pre-training the language encoder [51] to produce a 384-dimensional feature, which is subsequently projected to 512 dimensions through an MLP. Visual observations of size $480\times 640\times 3$ are encoded using ResNet18 [53], with a FiLM conditional layer [54] applied to each view encoding, ensuring robust task activation performance in multi-task scenarios[9]. The visual observations are finally transformed into a $(300\times n)\times 512$ feature sequence, where $n$ denotes the number of views used. The goal image $g_{i}$ is encoded by the pre-trained ResNet18, producing a $300\times 512$ feature, and remains fixed during training without parameter updates. The latent variable $z$ is obtained with a 4-layer transformer encoder and projected to 512 dimensions. Proprioceptive input $j_{t}$ is projected to 512 dimensions through an MLP. The CVAE decoder consists of a 4-layer transformer encoder and a 7-layer transformer decoder. The input feature dimensions for the transformer encoder are $(303+300\times n)\times 512$. The transformer encoder fuses features from different modalities, and the decoder outputs the predicted action chunk $\hat{a}_{t:t+k}$ and $k$ envisioned foresight scenes $\hat{f}_{t:t+k}$.
The FoAM training process, which incorporates the FA, is outlined in Algorithm 1. During training, we use L1 loss to compute the action loss $\mathcal{L}_{\text{action}}$ and Huber loss to compute the foresight loss $\mathcal{L}_{\text{foresight}}$, along with a KL divergence term $\mathcal{L}_{\text{reg}}$ regularizing the CVAE encoder. These losses are weighted by $\alpha$, $\beta$, and $\gamma$; throughout our experiments, the weight values set to $1$, $2$, and $10$, respectively.
During inference, the FA module is discarded, and the policy at this stage is represented by $\pi_{\theta}(\hat{a}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$. Based on the current observations and goals, the action chunk $c_{t}=\hat{a}_{t:t+k}$ is predicted. Following prior action chunk-based policies [9, 20, 10], we apply exponential temporal aggregation to produce smoother motion trajectories. Unlike previous work, we introduce the hyperparameter $r$ to eliminate the equality constraint on chunk size $k$ and the temporal aggregation range $r$. This is particularly crucial for deploying FoAM in real-world scenarios, as it allows flexibly adjusting the smoothing range according to the characteristics of different tasks, and optimizing task performance. The inference code is shown in Algorithm 2. There are approximately $160M$ parameters in the training and around $80M$ in the inference.
@@ III-D FoAM Policy Implementation
0:  Expert demo $\mathcal{D}$, maximum episode length $T$, chunk size $k$ ($k\approx T$), and loss weights $\alpha$, $\beta$, $\gamma$

### v2

@@ Abstract
Multi-task imitation learning (MTIL) has shown significant potential in robotic manipulation by enabling agents to perform various tasks using a single policy. This simplifies the policy deployment and enhances the agent’s adaptability across different scenarios. However, key challenges remain, such as maintaining action reliability (e.g., avoiding abnormal action sequences that deviate from nominal task trajectories) and generalizing to unseen tasks with a few expert demonstrations. To address these challenges, we introduce the Foresight-Augmented Manipulation Policy (FoAM), a novel MTIL policy that pioneers the use of multi-modal goal condition as input and introduces a foresight augmentation in addition to the general action reconstruction. FoAM enables the agent to reason about the visual consequences (states) of its actions and learn more expressive embedding that captures nuanced task variations. Extensive experiments on over 100 tasks in simulation and real-world settings demonstrate that FoAM significantly enhances MTIL policy performance, outperforming state-of-the-art baselines by up to 41% in success rate. Meanwhile, we released our simulation suites, including a total of 10 scenarios and over 80 challenging tasks designed for manipulation policy training and evaluation. See the project homepage https://projFoAM.github.io/ for project details.
@@ I Introduction
One of the main goals of robot learning is to develop a general-purpose agent capable of performing various tasks based on user commands. Multi-task imitation learning (MTIL) is a key approach that enables agents to learn multiple skills from expert demonstrations and trains an efficient policy, eliminating the need for complex, hardcoded solutions or reward functions. However, suboptimal policies and how to obtain generalization with limited training data are key issues that need to be addressed in MTIL policy development. This requires the novel MTIL framework to enable the agent to learn task-independent skills and more expressive information from demonstrations, capturing task-specific details to ensure reliable execution of individual tasks and generalizing to unseen tasks and scenarios [1, 2, 3].
@@
Previous research has shown that task adaptation in MTIL can be achieved by incorporating goal conditions into multi-task policy training [4, 5, 6, 7, 8, 9, 10, 11]. However, the low reliable execution of multi-task policies remains a challenge. Existing MTIL policies that align robotic actions with expert actions based on goal conditions often fail to reason about these distribution differences in the demonstrations of various tasks, severely impacting the agents’ performance on individual tasks. Meanwhile, most MTIL policies adopt uni-modal goal conditions [10, 12, 13, 14, 15, 16, 17, 9], which results in some inherent limitations. For example, with limited demonstrations and without data augmentation, policies conditioned on task instructions have difficulty generalizing to unseen tasks (discuss in Section IV-B). While policies conditioned on goal images offer fine-grained guidance and even demonstrate striking zero-shot capabilities [13], they encounter ambiguities during task execution. Ambiguity means that the same goal image may exist for different tasks in the same scenario. For instance, when the robot was tasked with placing an object into a multi-layer locker, the resulting goal image was ambiguous due to closedness. The manipulated object in the initial image disappeared in the goal image, making it impossible to determine which specific layer the object was placed in just based on the goal image (see Figure 4).
@@
In this paper, we introduce the Foresight-Augmented Manipulation Policy (FoAM), a novel MTIL policy designed to enhance the task performance of agents while addressing the limitations of uni-modal goal-conditioned policy. This approach is inspired by the perception ability of humans when performing tasks. When humans receive task instructions, they can easily foresee the state when completing the task in mind, and guide task execution based on this foresight and real-time actions until the expected results are achieved [18]. Similarly, FoAM addresses the limitations of uni-modal goal-conditioned MTIL policies by utilizing a multi-modal goal condition (goal image and task instruction). It also innovatively attempts to leverage a fine-tuned vision-language model (VLM) [19] to autonomously generate goal images. During training, we apply an action loss to refine the policy’s behavior and novelly introduce a foresight loss to control the consequences of its actions. This allows the agent to reason about its action across diverse tasks, and handle the ambiguities and variations in expert demonstration data, leading to more forward-looking and precise action during inference. FoAM demonstrates significant effectiveness with evaluations across more than 100 tasks in both simulation and the real world. It outperforms state-of-the-art baselines, achieving an increase in success rate by up to 41% in success rate. Our main contributions are summarized as follows:
  • We employ multi-modal goal condition to address the limitations of uni-modal goal-conditioned MTIL policies. FoAM demonstrates the capability to generalize to unseen tasks with limited expert demonstrations while ensuring efficient execution in ambiguous scenarios. Additionally, we attempt to involve a Vision-Language Model (VLM) in FoAM, which enables the agent to autonomously acquire goal images.
  • We propose foresight augmentation, a novel module designed to enable agents to learn more expressive embedding and enhance task performance by aligning task instructions with the consequences of their actions.
  • We have open-sourced the first simulated dual-arm system with rich tasks (over 80) and scenarios (over 10). This system replicates the widely-used UR3e robot in MuJoCo [20] and will be continuously maintained. It serves as a valuable simulation tool for developing robotic manipulation policies, such as MTIL and Sim2Real reinforcement learning policies [21].
@@ II Related Work
Goal-conditioned Learning for Robotic Manipulation. In recent years, significant progress has been made in single-task learning policies [22, 23, 24, 25, 26]. However, to enable a wider adaptability, intelligent robots must be equipped with the ability to conduct diverse tasks and complete them effectively. Among current MTIL approaches, language-conditioned policies utilize large-scale datasets to achieve task generalization, or apply data augmentation techniques, such as vision generation models, to modify backgrounds and manipulated objects, enabling generalization across more tasks and scenarios with limited training data [6, 27, 7, 9, 10, 28, 29, 8, 30, 31, 32, 33]. Despite the promising initial success, we found that language-conditioned policies often struggle with unseen tasks without sufficient data or extensive data augmentation. In parallel, some policies have introduced the goal image as task condition [12, 10, 13, 34, 35, 36]. Compared to language inputs, images could provide more expressive representation at the pixel level, enabling stronger generalization capabilities, and even allowing agents to perform zero-shot tasks [13]. However, goal images are susceptible to scenario ambiguity, where visually identical images can be produced by different tasks, causing actions inconsistent with intention.
@@
Collecting goal images requires human involvement, such as hand-drawing sketch [12], which reduces the autonomy of the agent. Recent work has explored multi-modal goal conditions to enhance agent’s task performance [15, 14, 37, 38, 36]. These policies extract workspaces and manipulated objects from both task instruction and initial observation, then compose a multi-modal embedding based on predefined templates. In contrast, FoAM leverages a vision-language model to generate a goal image with semantic information. The generated goal image and task instruction are directly processed to infer actions and predict their consequences. By doing so, our method addresses the limitations of uni-modal goal-conditioned policies that rely solely on language instructions, goal images, or predefined templates.
@@
Agents with Vision Language Models. In recent years, Vision-Language Models (VLMs) have been introduced to robotics [32, 39, 40, 41, 42, 43, 38, 44, 45], enabling more complex visual reasoning and multi-modal application. Meanwhile, in the community of image editing, VLMs also demonstrated the ability to understand language instructions, editing real-world images and producing highly realistic visual effects[46, 35]. Works[47, 48, 49] further validate that edited images can be interpreted by robotic agents, where the generated goal images are directly used as a goal condition. In contrast, we integrate VLMs seamlessly into our framework not only during the inference stage, but also in policy training. The generated goal images serve as the “labels” to compute the reconstruction loss for the policy (Section III-C), coupling the action learning and action result prediction.
@@ III METHOD
We seek to develop an MTIL policy that can reliably generate actions consistent with semantics while effectively generalizing to unseen tasks with a small amount of training data, enabling agents to complete various tasks more efficiently and accurately. To achieve this, we introduce FoAM, a novel multi-modal goal-conditioned policy. In the following sections, we will provide an overview of the FoAM in III-A, detail the fine-tuning of a cutting-edge visual-language model to generate goal images for FoAM in III-B, propose the foresight augmentation in III-C, and introduce the implementation details of FoAM in III-D.
Figure 1: Training and inference pipelines of FoAM. The input terms remain consistent throughout both training and inference. During training, actions and their corresponding consequences are predicted, with the predicted consequences and actions being aligned with the input goal image and expert actions, respectively, to update the parameters of FoAM. During inference, the trained policy is used solely to predict the action $\hat{a}_{t}$.
@@ III-A Pipeline Overview
The pipeline of FoAM is illustrated in Figure 1. FoAM is a transformer-based [50] policy that inherits the architecture of the prior work [22] and is trained as a conditional variational autoencoder (CVAE) [51, 52]. The process begins with the user providing a task instruction $g_{l}$ to the agent, which is then fed into FoAM as a language goal condition through a pre-trained text encoder [53]. Concurrently, the task instruction, such as Put eggplant into the green bowl, is input into the Goal Imagination Module along with the initial observation $o_{i}$. The generated goal image or human-collected real goal image $g_{i}$ serves as image goal condition. By integrating the agent’s proprioception $j_{t}$, visual observations $o_{t}$, and the latent style variable $z$ encoded from all episodes, the multi-modal goal-conditioned policy $\pi_{\theta}(\hat{a}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$ is reconstructed by the coupling of foresight and action loss. During inference, action chunks are predicted and smooth actions are produced through temporal aggregation [22].
@@ III-B Fine-tuned Goal Imagination Module
The acquisition of goal images often requires human participation, such as hand-drawing sketch [12] or collecting real goal images [10] before inference. In order to improve the efficiency of the agent’s task execution, we embed a goal imagination module in FoAM to autonomously generate goal images. In this work, we chose InstructPix2Pix (Ip2p)[19] as the goal imagination module, leveraging a dataset of over 20,000 training pairs. Of these, 16,000 pairs were obtained from robot expert demonstrations provided by RT-1[27, 6, 7]. The first and last frames of these demonstrations were used as the original and edited images, respectively, with the corresponding task name serving as the instruction. Since RT-1 demonstrations exhibited perturbations in the final frames due to robot arm movements, we implemented a rigorous data-cleaning procedure to remove noise and ensure high-quality training data. Additionally, we incorporated over 4,000 data pairs from our own simulation and real-world datasets (see Section IV-A for details).
We fine-tuned Ip2p for 500 epochs on a single NVIDIA H100 GPU, a process that required approximately 3 days. During the inference stage, with the model weights pre-loaded, generating one goal image of size $480\times 640\times 3$ took about 4 seconds. Figure 2 presents some inference demonstrations captured during VLM-FoAM joint inference experiments (Section IV-C). They highlight the model’s ability to generate realistic and semantically consistent visuals based on the given initial observations and task instructions.
@@
Figure 2: Inference demonstrations of the fine-tuned goal imagination module. The leftmost image illustrates the initial observation, while the next four images represent the edited goal images generated based on the initial observation and the task instruction provided at the top. Please visit the project homepage for more examples.
III-C Foresight Augmentation
@@
Humans possess exceptional perception abilities for understanding and interacting with external events. When performing a task, they can easily foresee the goal state of a scenario before execution in mind and use it as a guide. Inspired by this capability, we developed a key module called Foresight Augmentation (FA) to equip agents with a similar perception mechanism. FA enables the agent to simultaneously comprehend both its actions and the subsequent consequences these actions will produce. By integrating FA into the agent’s decision-making process, we significantly enhance its overall performance in tasks execution (see Section IV-B for details).
@@
We train FoAM as a CVAE, utilizing an encoder similar to those in [22, 9, 54, 24] to generate the latent variable $z\sim q_{\phi}(z|a_{t:t+k},{j}_{t})$. The decoder is defined as policy $\pi_{\theta}(\hat{a}_{t:t+k},\hat{f}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$, which predicts a $k\times n$-dimensional action chunk $\hat{a}_{t:t+k}$ and foresight sequence $\hat{f}_{t:t+k}$ based on real-time observation and conditions, where $k$ represents the hyperparameter chunk size during training and $n$ denotes the dimension of the agent’s action space. To enhance the agent’s ability to interpret and respond to the dynamic work scenario, we strategically increase the value of $k$, thereby expanding the agent’s foresight horizon. For example, since each episode in the multi-task expert demonstration dataset $\mathcal{D}$ often have different time steps, we define the maximum time step of all episodes as $T$ and set $k$ to a value close to $T$. During training, while FoAM predicts the action chunk $\hat{a}_{t:t+k}$, FA generates $k$ potential foresight $\hat{f}_{t:t+k}$ and selects the frame $\hat{g}_{i}=\hat{f}_{t:t+k}[k-t]$ that is temporally consistent with the goal image $g_{i}$. This selected frame $\hat{g}_{i}$ is then aligned with the goal image $g_{i}$ using the foresight loss $L_{\text{foresight}}$ calculated by the Huber Loss, which is coupled with the action loss $L_{\text{action}}$ as the reconstruction loss $L_{\text{recon}}$ of the policy to play a role in updating the policy’s parameters (Algorithm 1). This process emulates the strong perception abilities humans exhibit when performing tasks, and we demonstrated in experiments that the FA significantly enhances the agent’s task performance (Section IV-B).
@@
0:  Expert demo $\mathcal{D}$, maximum episode time step $T$, chunk size $k$ ($k\approx T$), and loss weights $\alpha$, $\beta$, $\gamma$
@@
III-D FoAM Policy Implementation
FoAM is designed as a transformer-based policy with sufficient capacity to predict specific sequences by effectively integrating sequence information from the inputs. FoAM is implemented using an ACT-like architecture [22] with the CVAE framework. The language-conditioned embedding is obtained by the pre-trained language encoder [53] to produce a 384-dimensional feature, which is subsequently projected to 512 dimensions through an MLP. ResNet18 [55] with FiLM conditional layer [56] is used to encode visual observations of size $480\times 640\times 3$ and embed task instructions into images, ensuring robust task performance in multiple scenarios [9]. The visual observations are finally transformed into a $(300\times n)\times 512$ feature sequence, where $n$ denotes the number of used viewpoints. The goal image $g_{i}$ is encoded by the pre-trained ResNet18, producing a $300\times 512$ feature, and remains fixed during training without parameter updates. The latent variable $z$ is obtained with a 4-layer transformer encoder and projected to 512 dimensions. Proprioceptive input $j_{t}$ is projected to 512 dimensions through an MLP. The CVAE decoder consists of a 4-layer transformer encoder and a 7-layer transformer decoder. The input feature dimensions for the transformer encoder are $(303+300\times n)\times 512$. The encoder fuses features from different modalities, and the decoder predicts the action chunk $\hat{a}_{t:t+k}$, while generating $k$ foresight $\hat{f}_{t:t+k}$ (each $300\times 512$) through a second-dimensional full connection layer.
The FoAM training process, which incorporates the FA, is outlined in Algorithm 1. During training, we use L1 Loss and Huber Loss to compute the action loss $\mathcal{L}_{\text{action}}$ and foresight loss $\mathcal{L}_{\text{foresight}}$ respectively, along with a KL divergence term $\mathcal{L}_{\text{reg}}$ regularizing the CVAE encoder. These losses are weighted by $\alpha$, $\beta$, and $\gamma$. In our experiments, the weight values are set to $1$, $2$, and $10$, respectively.
During inference, the FA module is discarded. The implemented policy is represented by $\pi_{\theta}(\hat{a}_{t:t+k}|o_{t},j_{t},g_{i},g_{l},z)$. Based on the current observations and goals, the action chunk $c_{t}=\hat{a}_{t:t+k}$ is predicted. Following prior action chunk-based policies [9, 22, 10], we apply exponential temporal aggregation to produce smooth action trajectories. Unlike previous work, we introduce the hyperparameter temporal aggregation range $r$ as the actual chunk size during inference, which eliminates the equality constraint on chunk size $k$ during training and inference. This is particularly crucial for deploying FoAM, as it allows flexibly adjusting the aggregation range according to the user’s intentions and the characteristics of different tasks, and optimizing task performance. During experiments, we observed that when $r$ is large ($>100$), the agent can perform tasks more efficient, and when $r$ is relatively small ($50\sim 80$), the agent has better reactiveness and ability to resist external disturbance (Section IV-D). The inference code is shown in Algorithm 2. There are approximately $160M$ parameters in the training and around $80M$ in the inference.